OpenStudy (anonymous):

Solve the equation for y:) 2x+7y=14

OpenStudy (littlenugget):

What part do you need help in?

OpenStudy (littlenugget):

;)

OpenStudy (anonymous):

Dividingb it by 7

OpenStudy (anonymous):

\[2x+7y=14\] 7y=-2x+14 y=-2/7x+2

OpenStudy (littlenugget):

ok, so basically: 2x+7y=14 -2x -2x -------------- 7y=-2x+14 So this is where you got to right? when you divide by seven, you are just going to divide everything by 7: like -2x/7 +14/7= y So what would you get?

OpenStudy (littlenugget):

when you simplify?

OpenStudy (littlenugget):

You should get what @jerryjjoby put :)

OpenStudy (anonymous):

Thanks a lot!!!:)

OpenStudy (littlenugget):

no prob :)

OpenStudy (anonymous):

Glad to assist.